Please write out the complete mechanism using curved arrow formalism, showing each intermediate expected along the reaction pathway for synthesis of 2-butoxynaphthalene. Ethanol (20 mL), NaOH (554 mg), 2-naphthol (997 mg) and 1-bromobutane (1 mL) were used in this experiment. Also, could you please show step-by-step calculations of limiting reactant and theoretical yield?
Answer #1

Synthesis of 2-butoxynaphthalene

Mechanism:

Na OH + N + H2O 2-naphthol naphthoxide ion 00 2-butoxynaphthalene

No. of moles of NaOH =mass/molar mass = 554/40 = 13.85 mmol

No. of moles of 2-naphthol = 997/ 144.2 = 6.91 mmol = 0.00691 moles

No. of moles of 1-bromobutane = mass/137

mass = volume x density = 1 x 1.27 = 1.27 g

no.of moles of 1-bromobutane = 1.27/137 = 9.27 mmol

So here the limiting reagent is 2- naphthol (since the stochiometry of the three reagent in the balanced chemical equation is 1:1:1 )

Theoretical yield = no.of moles of limiting reagent x molar mass of 2-butoxynaphthalene

   = 0.00691 x 200.28 = 1.3839 g